Sid can paint a house by himself in 10 hours and Sid and Bill can paint a house together in 4 hours. How long does it take Bill to paint a house by himself?
Answer by ptaylor(2198)

Let x=amount of time it takes Bill to paint the house by himself
Then Bill paints at the rate of 1/x house per hour
Sid paints at the rate of 1/10 house per hour
Sid and Bill together paints at the rate of 1/4 house per hour
soooooo:
1/10+1/x=1/4 multiply each term by 40x
4x+40=10x
6x=40
x=6 two thirds hours-------time it takes Bill to paint the house by himself
CK
1/10+1/(20/3)=1/4
1/10+3/20=1/4
2/20+3/20=1/4
5/20=1/4
1/4=1/4
Hope this helps---ptaylor